# Towards Precise 3D Quantum Control of a Levitated Dipolar Scatterer using Spatial Mode Decomposition

### v1

@@ Abstract
An essential requirement for universal quantum control of the centre-of-mass motion of levitated objects is the development of a precise readout of all three translational degrees of freedom. Improving that precision presents one key challenge: collecting all the information on the object’s position, encoded in the scattered light, equivalent to minimising the measurement imprecision. Here, we propose a new detection technique based on spatial mode decomposition, which addresses this problem using a simple integrated setup, where all of the light back-scattered from a levitated nanoparticle couples into a spatial mode sorter. We observe that each mode of the sorter pairs predominantly to the in-elastically scattered field generated by the object’s motion along a particular spatial axis. This results in each translational degree of freedom being selectively encoded in the amplitude of orthogonal information channels. Using this approach, we report measurement efficiencies ($\eta_{\mathrm{tot}}^{x}$, $\eta_{\mathrm{tot}}^{y}$, $\eta_{\mathrm{tot}}^{z}$) = (0.17, 0.15, 0.30), implying that our technique is capable of reaching the 3D quantum ground state.
@@
Levitated optomechanical systems have recently entered the quantum regime, with several demonstrations having now achieved cooling to the centre-of-mass (COM) motional ground state [1, 2, 3, 4]. Such platforms present an exciting avenue for designing tests of fundamental physics [5, 6, 7, 8], and for matter-wave interferometry experiments with mesoscopic objects [9] several orders of magnitude larger than the current state-of-the-art [10, 11]. To achieve ground-state cooling, research efforts can be separated into two distinct approaches: passive cooling using coherent scattering in a cavity [1, 12, 13], and active cooling with measurement-based control techniques [2, 3]. In particular, the latter actively controls an object’s mechanical motion by applying a position or velocity-dependent feedback force. Hence, measurement-based control techniques are highly dependent on the efficiency with which we can collect information on the system’s dynamics.
So far, conventional interferometric techniques have only reached quantum enabling measurement efficiencies for the motion parallel to the propagation axis of the optical tweezer, and often at the detriment of the transverse translational degrees of freedom (DOFs). In an attempt to reconcile this, research efforts have focused on developing more sophisticated detection systems, with a particular emphasis on investigating the spatial distribution of information in the scattered power, referred to as the information radiation patterns [14, 15].
@@
Building on these efforts, we present a novel approach by decomposing the in-elastically scattered field in the appropriate basis of spatial modes. This allows us to isolate contributions from all three translational degrees of freedom. As a convenient demonstration of this technique, we choose the linearly polarised (LP) mode basis of an optical fiber, encoding information on each translational DOF in the amplitude of a unique mode. Our demonstration relies on a simple and integrated experimental setup based on telecommunication grade, commercially available components. Finally we perform parametric feedback cooling, and show that our detection technique allows us to reach the measurement efficiencies necessary for realizing the 3D ground state of a levitated particle.
I Sorting the Scattered Light
@@
Figure 1: (a) The operating principle of our displacement sensing technique using a spatial mode sorter. In the focal plane of the collection lens, labelled [1], the scattered field and reference field are coupled into the few-mode optical fiber input of the device. The in-elastically scattered field component $\bm{E}^{(q)}(x,y)$ carrying information on the nanoparticle’s motion along the $\bm{\hat{e}}_{q}$ axis couples predominantly to a single LP mode. In the output plane, labelled [2], information on each translational degree of freedom $q(t)$ is stored in a unique channel of the sorter. (b) The output of a $\lambda=1550$ nm laser is controlled by an acousto-optic modulator (AOM) and coupled into the LP01 channel of a space-division de-multiplexing photonic chip, producing a $100$ mW optical tweezer at the focus of a high numerical aperture parabolic mirror. The back-scattered light from a levitated nanoparticle is then coupled back into the photonic chip, allowing us to measure each translational degree of freedom independently.
@@
We are concerned with the case of a sub-wavelength isotropic dipolar scatterer, referred to as a nanoparticle, confined to the focus of a linearly polarised trapping beam. In particular, we model the geometry shown in Fig. 1(a), where the trapping beam is focused by a high-numerical aperture (NA) parabolic mirror. This implementation presents a key advantage over traditional setups [16], in that it collects all of the back-scattered light, which is collimated and subsequently coupled into a few-mode optical fiber in the focal plane of a collection lens. At the output of the few-mode fiber, we measure the recovered power to extract information on the nanoparticle’s real-time position.
@@
In the limit of small displacements, we find that we can write the scattered field in the focal plane of the collection lens as a sum of elastically and in-elastically scattered light (see Supplementary Material). In particular, we find that only the in-elastically scattered light is modulated by the nanoparticle’s real-time position and thus is the only one to contribute meaningfully to the measurement result. That is,
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))=\bm{E}_{0}(x,y)+\sum_{q=x,y,z}\frac{q(t)}{w_{0}}\bm{E}^{(q)}(x,y)~{},$ (1)
@@
where $w_{0}$ refers to the trapping beam waist at the focus of the parabolic mirror, and $q(t)$ refers to the nanoparticle’s position along the $\bm{\hat{e}}_{q}$ axis. In addition to these terms, we must also consider the un-scattered and diverging part of the trapping beam, $\bm{E}_{\mathrm{ref}}(x,y)$, which functions as the local oscillator (LO) in conventional interferometry-based detection systems. In the focal plane of the collection lens, this field is well-approximated by a plane wave and can usually be tuned such that ${|\bm{E}_{\mathrm{ref}}(x,y)|\geq|\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))|}$.
@@
From Eq. (1), it is clear that the spatial mode of the in-elastically scattered field is highly dependent on the displacement axis of the nanoparticle, $\hat{\bm{e}}_{q}$. In particular, displacements parallel to the trapping axis ($\bm{\hat{e}}_{z}$) produce the spatial-intensity pattern shown in Fig. 1(a), which we note is reminiscent of the LP01 mode supported in optical fibers. Conversely, the patterns produced for displacements along $\bm{\hat{e}}_{x}$ and $\bm{\hat{e}}_{y}$ are reminiscent of the LP11a and LP11b modes, respectively. Inspired by this observation, we have opted to couple the scattered field into a space-division de-multiplexing photonic chip, which functions as a spatial mode sorter, to isolate contributions from each translational degree of freedom (DOF) in a unique LP mode.
II Experimental Setup
@@
We optically levitate spherical SiO2 nanoparticles of diameter $310$ nm, using the experimental setup illustrated in Fig. 1(b). We load the trap at ambient pressure using an ultrasonic nebulizer and a dilute solution of sonicated SiO2 suspended in ethanol. The trapping beam is tuned to a wavelength of $\lambda=1550$ nm and set to a power of approximately $100$ mW. Upon loading the optical trap, the system is decompressed in a two-stage pumping process to a vacuum on the order of $5\times 10^{-5}$ mbar.
In the focal plane of the collection lens, the scattered field is coupled into the spatial mode sorter [17]. The input is a few-mode, graded-index optical fiber with a core size of approximately 8 µm, capable of supporting the LP01, LP11a and LP11b modes. Once excited, the power in each of these modes is transferred to a unique single-mode output fiber using a tapered mode-selective coupler [18, 19]. The power in each output channel is sent to a low-noise photodiode, with a typical set of detected signals shown in Fig. (2). For a horizontally-polarized trapping beam, we observe oscillation frequencies ${(\Omega_{x},\Omega_{y},\Omega_{z})~{}/~{}2\pi\approx(106,42,54)}$ kHz.
For parametric feedback cooling, the signal of each photodiode is sent to an FPGA board, which functions as a digital phase-locked loop (PLL), with the recovered signal filtered and phase-shifted to generate a feedback signal at twice the natural oscillation frequency, $2\Omega_{q}$. These feedback signals are used to modulate the power of the trapping beam via an acousto-optic modulator (AOM).
III Results
We find that each output channel in Fig. 2(a) contains information on primarily a single translational DOF, equivalent to the nanoparticle’s motion along a particular spatial axis. This information is encoded in the power spectral density (PSD) $S_{qq}(\Omega)$ for $q=x,y$ and $z$. We observe good extinction ratios, with the dominant spectral feature in each channel being at least two orders of magnitude larger than all others. To quantify this selectivity, we have numerically modelled the coupling efficiency of the in-elastically scattered field to each LP mode, defined
@@ III Results
where $\textbf{LP}_{\ell,m}(x,y)$ refers to the electric field of the LPℓ,m mode, and integration is performed over the end-face of the input optical fiber (see Supplementary Material). For the particular experimental setup used here, this produces the values reported in Fig. 3 below, where we have also opted to include the coupling efficiencies for the elastically scattered and LO fields.
@@
Figure 3: The coupling efficiency of each electric field component to each LP mode supported by the spatial mode sorter, in the focal plane of the collections lens. We find that the in-elastically scattered field components, $\bm{E}^{(q)}(x,y)$, selectively couple to specific LP modes. Note that the LP11a/b modes are effectively de-coupled from the elastically scattered and LO fields, $\bm{E}_{0}(x,y)$ and $\bm{E}_{\mathrm{ref}}(x,y)$, respectively.
Interestingly, the elastically scattered and LO fields are effectively de-coupled from the LP11a/b modes on account of their spatial asymmetry. This de-coupling would typically be disadvantageous in conventional detection systems, which rely on the LO to extract information from the phase of the scattered light. However, this is not the case with our approach. Indeed, we find that the measurement imprecision is proportional only to the coupling efficiency of the in-elastically scattered field (see Supplementary Material). This relaxation of the strict requirements imposed on conventional detection systems suggests that our technique remains near optimal despite filtering out significant amounts of the incident light. This result is reflected in the signal-to-noise ratios observed in Fig. (2b), which we note are unusually high for the transverse DOFs, with an optimal value of ${\mathrm{SNR}_{x}^{\mathrm{LP_{11a}}}\approx 110}$ dB above the shot noise floor.
@@
Finally, note that the LO field usually also serves to offset the recovered signal, ensuring it sits above the system’s technical noise. Since the LO does not couple into the LP11a/b modes to provide this offset, the power in these channels is much weaker than in the LP01 mode ($<10\%$). Consequently, if one were to measure every LP mode simultaneously with a single detector, $S_{zz}(\Omega)$ would dominate the measurement result, and the transverse DOFs would be difficult to discern. This is why a spatial mode sorter must be used, as opposed to a simple few-mode fiber configuration, and explains why the benefits of this approach have yet to be observed in previous experiments.
@@
As a proof-of-principle demonstration of our detection technique, we use parametric feedback cooling to control the COM motion of a levitated nanoparticle. The temperature $T_{q}$ achieved along the $q$-axis is determined by fitting the PSD measured at the appropriate output channel of the photonic chip (see Supplementary Material). To capture the actual dynamics of the system, the collected signals have been averaged over many oscillatory periods. We find that the motion of the nanoparticle can be cooled from thermal equilibrium ($300$ K) to a minimum temperature of ${(T_{x},T_{y},T_{z})=(100,7,20)}$ mK, respectively. Owing to record high measurement efficiencies, the minimum temperatures achieved here are an order of magnitude below what has been reported in the literature at pressures of approximately $5\times 10^{-5}$ mbar [16], the current technical limitation of our vacuum chamber. Indeed, we find that our observed values are approaching the theoretically predicted minimum values at this pressure (see Supplementary Material).
@@
We note that $T_{y}$ is the lowest observed temperature despite having the worst reported measurement efficiency. We believe this is because $\Omega_{x}\approx 2\Omega_{z}$, such that the feedback signal generated for the longitudinal DOF near-resonantly drives the nanoparticle’s motion along the $\bm{\hat{e}}_{x}$ axis. This can be avoided by varying the ellipticity of the polarisation of the trapping beam, which tunes the transverse oscillation frequencies [20]. However, this also leads to modifications in the spatial profile of the scattered field to the detriment of the demultiplexing of the transverse DOFs.
IV Discussion
Measurement-based quantum control of an object’s mechanical motion is inherently dependent on measurement efficiency, which represents the amount of useful information that is not lost to the environment. To understand this, note that optical measurements necessarily disturb the system’s natural evolution due to back-action. However, efficient measurements maximise the information gained per disturbance and can generate real-time feedback, which not only controls the system’s mechanical motion but also cancels the effects of said back-action.
In clamped optomechanical systems, the measurement efficiency can be made arbitrarily high [21], but the intrinsic cross-coupling between different mechanical modes (i.e., background mechanical modes) has been a persistent issue for multidimensional quantum control [22]. This is less of a concern in levitated optomechanics, where the orthogonality of each mechanical DOF means it is relatively straightforward to isolate contributions from a given mode, provided there is no residual libration or rotation. However, until the present demonstration, realising the measurement efficiencies required to precisely control one mechanical DOF has come to the detriment of capturing the full 3D motion of a levitated mechanical oscillator.
@@ IV Discussion
Coupling Efficiency $\begin{cases}0.66&,~{}x(t)\to\mathrm{LP}_{11\mathrm{a}}\\ 0.60&,~{}y(t)\to\mathrm{LP}_{11\mathrm{b}}\\ 0.67&,~{}z(t)\to\mathrm{LP}_{01}\\ \end{cases}$
@@
Measurement Efficiency ($\eta^{q}_{\mathrm{tot}}$) $\begin{cases}0.17&,~{}x(t)\to\mathrm{LP}_{11\mathrm{a}}\\ 0.15&,~{}y(t)\to\mathrm{LP}_{11\mathrm{b}}\\ 0.30&,~{}z(t)\to\mathrm{LP}_{01}\end{cases}$
@@
Table 1: Total efficiencies for the setup considered here. For a discussion on how each term has been determined, refer to the Supplementary Material.
We outline the information efficiencies of our system in Table (1), where we report record measurement efficiencies for the transverse DOFs. Importantly, this is not at the expense of the longitudinal DOF, for which we achieve a value comparable to previous experimental demonstrations of 1D ground state cooling [3, 2].
@@
From these values, we can estimate the minimum achievable phonon occupancy for the nanoparticle’s COM motion along each spatial axis [23, 21, 3] in the limit of low pressures, where environmental information loss from random collisions with background gas molecules is negligible (${\approx 1\times 10^{-8}}$ mbar) [24]. Under these conditions, with an optimised feedback circuit, we expect to reach phonon occupancies of ${\bar{n}_{\mathrm{min}}^{q}=(1/\sqrt{\eta^{{}_{q}}_{{}^{\mathrm{tot}}}}-1)/2}$. For the setup reported here, this is equivalent to ${(\bar{n}_{\mathrm{min}}^{x},\bar{n}_{\mathrm{min}}^{y},\bar{n}_{\mathrm{min}}^{z})=(0.71,0.79,0.41)}$. Hence, this is the first demonstration of a detection system capable of realising the 3D motional ground state of a levitated nanoparticle using measurement-based control techniques.
V Summary
@@
In summary, we have realized a novel approach to measuring the centre-of-mass motion of a levitated dipolar scatterer, wherein the scattered field couples into an on-chip spatial mode sorter. By decomposing the in-elastically scattered contributions to this field in the LP mode basis of an optical fiber, we can isolate contributions from each translational degree of freedom in the amplitude of orthogonal information channels. Thanks to this technique, we are able to report record measurement efficiencies for the transverse degrees of freedom, which suggests that our technique is capable of realizing the 3D motional quantum ground state. Importantly, we believe that our technique can be extended to higher-order modes in the LP basis and generalized to other families such as Hermite-Gaussian (HG) modes, orbital angular momentum (OAM) modes, and Laguerre-Gaussian (LG) modes. This opens an exciting avenue to tracking the entire 6D motion of levitated objects with arbitrarily high collection efficiencies. Alternatively, one could harness the toolbox of structured light to match the in-elastically scattered field of the levitated object perfectly.
It is worth acknowledging that although the present demonstration focuses on the field of levitated optomechanics, it has the potential to be employed with other trapped microscopic objects such as atoms, ions or molecules. Finally, the ability of the spatial mode sorter to isolate particular motional degrees of freedom without measurement is an essential step towards all-optical coherent feedback control [25] of a levitated particle.

### v2

# Three-Dimensional and Selective Displacement Sensing of a Levitated Nanoparticle via Spatial Mode Decomposition

@@ Abstract
We propose and experimentally demonstrate a novel detection method that significantly improves the precision of real-time measurement of the three-dimensional displacement of a levitated dipolar scatterer. Our technique relies on the spatial mode decomposition of the light scattered by the levitated object, allowing us to simultaneously and selectively extract the position information of all translational degrees of freedom with minimal losses. To this end, we collect all the light back-scattered from a levitated nanoparticle using a parabolic mirror and couple it into a spatial mode sorter. The sorter effectively demultiplexes the information content of the scattered electric field, resulting in each of the nanoparticle’s translational degrees of freedom being selectively encoded in the amplitude of orthogonal optical modes. We report measurement efficiencies of ${(\eta_{{}^{\mathrm{tot}}}^{{}_{x}},\eta_{{}^{\mathrm{tot}}}^{{}_{y}},\eta_{{}^{\mathrm{tot}}}^{{}_{z}})=(0.14,0.16,0.32)}$ $>$ 1/9, which should enable the 3D motional quantum ground state of a levitated optomechanical system. Further, we believe this technique opens up the possibility to implement coherent feedback control of a levitated nanoparticle.
@@
Levitated optomechanical systems have recently entered the quantum regime, with several demonstrations having now achieved cooling to the center-of-mass (COM) motional quantum ground state [1, 2, 3, 4]. Such platforms present an exciting avenue for designing tests of fundamental physics [5, 6, 7, 8], and in particular for matter-wave interferometry experiments with mesoscopic objects [9] several orders of magnitude larger than the current state-of-the-art [10, 11]. A common prerequisite for most of these proposals is to prepare the levitated mechanical oscillator in its motional quantum ground state. To achieve ground-state cooling, research efforts can be separated into two distinct approaches: passive cooling using coherent scattering in a cavity [1, 12, 13], and active cooling with measurement-based control techniques [2, 3]. In particular, the latter actively controls an object’s mechanical motion by applying a position or velocity-dependent feedback force. Hence, measurement-based control techniques are highly dependent on the efficiency with which we can collect information on the system’s dynamics.
It was recently understood that information on the different motional degrees of freedom (DOFs) are encoded in specific spatial radiation patterns, so-called information radiation patterns [14, 15]. Because of this non-uniformity, it is a difficult task to construct a detection system capable of efficiently and simultaneously measuring each motional DOF of a levitated nanoparticle. In particular, this requires either a carefully engineered local oscillator to optimally interfere with the scattered light [16] or position-sensitive detectors (such as a quadrant photodiode) that are blind to specific DOFs. The latter approach would intrinsically require beam-splitting, introducing extensive optical losses. As a result, so far, conventional interferometric techniques have only been able to reach quantum-enabling measurement efficiencies for the motion parallel to the propagation axis of the optical tweezer [2, 3, 4], often at the detriment of the transverse translational DOFs.
@@
In this work, we present a novel approach to displacement sensing that relies on decomposing the light scattered from a levitated nanoparticle in the appropriate basis of spatial modes. This allows us to passively and efficiently isolate contributions from all three translational DOFs without the need for direct measurement. This kind of filtering, which utilises properties of the scattered electric field rather than the scattered power, preserves the coherence of the collected light. Consequently, it holds promise for use in all-optical coherent feedback control schemes, which avoid the decoherence and back-action associated with a quantum measurement [17, 18, 19]. As a convenient demonstration of our method, we choose the linearly polarised (LP) mode basis of an optical fiber, encoding information on each translational DOF in the amplitude of a unique LP mode. The experimental setup relies on simple, telecommunication grade and commercially available components, providing a path towards a compact and fully integrated levitated optomechanical platform. Finally, we perform parametric feedback cooling [20], achieving temperatures of ${(T_{x},T_{y},T_{z})=(7,100,20)}$ mK at pressures of $5\times 10^{-5}$ mbar, and show that our detection technique allows us to reach the measurement efficiencies necessary for realizing the 3D motional quantum ground state of a levitated nanoparticle.
@@
Figure 1: (a) The operating principle of our displacement sensing technique using a spatial mode sorter. In the focal plane of the collection lens, labelled [1], the scattered field and reference field are coupled into the few-mode optical fiber input of the device. In the panel [1] on the right, we show the dominant components of the inelastically scattered field $E_{x}^{(q)}(x,y)$, which carries information on the nanoparticle’s motion along the $\bm{\hat{e}}_{q}$ axis, and couples predominantly to a single LP mode. In the output plane, labelled [2], information on each translational degree of freedom $q(t)$ is stored in a unique channel of the sorter (see panel [2] on the right for the plots of the electric fields of the three fiber modes). (b) The output of a $\lambda=1550$ nm laser is controlled by an acousto-optic modulator (AOM) and coupled into the LP01 channel of a space-division demultiplexing photonic chip, producing a $100$ mW optical tweezer at the focus of a high numerical aperture parabolic mirror. The back-scattered light from a levitated nanoparticle is then coupled back into the photonic chip, allowing us to measure each translational degree of freedom independently.
@@
We consider the case of a sub-wavelength nanoparticle, which is well-approximated as an isotropic dipolar scatterer, confined to the focus of a linearly ($x$) polarised trapping beam. In particular, we model the geometry shown in Fig. 1(a), where the trapping beam is focused by a high-numerical aperture (NA) parabolic mirror. This implementation presents a key advantage over traditional setups [21], in that it collects all of the back-scattered light, which is collimated and subsequently coupled into a few-mode optical fiber in the focal plane of a collection lens. At the output of the few-mode fiber, we measure the recovered power to extract information on the nanoparticle’s real time position.
@@
In the limit of small displacements, we can write the scattered field in the focal plane of the collection lens as a sum of elastically and inelastically scattered light (see Supplementary Material, Appendix A). In particular, we find that only the inelastically scattered light is modulated by the nanoparticle’s real time position and thus is the only one to contribute meaningfully to the measurement result. To first order, we approximate the scattered field as
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))=\overbrace{\bm{E}_{0}(x,y)}^{\text{elastic}}+\sum_{q=x,y,z}\frac{q(t)}{w_{0}}\overbrace{\bm{E}^{(q)}(x,y)}^{\text{inelastic}}~{},$ (1)
@@
where $w_{0}$ refers to the trapping beam waist at the focus of the parabolic mirror, and $q(t)$ refers to the nanoparticle’s position along the $\bm{\hat{e}}_{q}$ axis. In addition to these terms, we must also consider the unscattered and diverging part of the trapping beam, $\bm{E}_{\mathrm{ref}}(x,y)$, which functions as the local oscillator (LO) in the simplest conventional interferometry-based detection systems. In the focal plane of the collection lens, this field is well-approximated by a plane wave and can usually be tuned such that ${|\bm{E}_{\mathrm{ref}}(x,y)|\gtrsim|\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))|}$.
@@
In Fig. 1(a) we compare the dominant polarisation component of the inelastically scattered field $\bm{E}^{(q)}(x,y)$ to the lowest-order LP modes. We find that displacements parallel to the trapping axis ($\bm{\hat{e}}_{z}$) produce an electric field profile reminiscent of the LP01 mode supported in optical fibers. Conversely, the fields produced for displacements along $\bm{\hat{e}}_{x}$ and $\bm{\hat{e}}_{y}$ are reminiscent of the LP11a and LP11b modes, respectively. Inspired by this observation, we have opted to couple the scattered field into a space-division demultiplexing photonic chip [22], which functions as a spatial mode sorter. As depicted schematically in Fig. 1(a), this allows us to isolate contributions from each translational DOF in amplitude modulations of a unique LP mode. Before quantifying this effect, however, let us first consider the experimental setup used to implement this selective measurement protocol.
@@
We optically levitate spherical SiO2 nanoparticles of diameter $310$ nm, using the experimental setup illustrated in Fig. 1(b). We load the trap at ambient pressure using an ultrasonic nebulizer and a dilute solution of sonicated SiO2 suspended in ethanol. The trapping beam is tuned to a wavelength of $\lambda=1550$ nm and set to a power of approximately $100$ mW. Upon loading the optical trap, the system is decompressed in a two stage pumping process to a vacuum on the order of $5\times 10^{-5}$ mbar.
In the focal plane of the collection lens, the scattered field is coupled into the spatial mode sorter [22]. The input is a few-mode, graded-index optical fiber with a core diameter of approximately 8 µm, supporting the LP01, LP11a and LP11b modes. Once excited, the electric field associated with each of these modes is transferred to a unique single-mode output fiber using a tapered mode-selective coupler [23, 24]. Each output channel is then sent to a low-noise photodiode, with a typical set of detected signals shown in Fig. 2. For an $x$-polarized trapping beam, we observe oscillation frequencies ${(\Omega_{x},\Omega_{y},\Omega_{z})~{}/~{}2\pi\approx(42,106,54)}$ kHz for each translational DOF. The other spurious spectral features in Fig. 2(a) are a consequence of the large nanoparticles used here, which experience coupling between different translational DOFs at room temperature ($300$ K) [25]. This leads to the generation of spectral features at frequencies given by various linear combinations of $\Omega_{x}$, $\Omega_{y}$, and $\Omega_{z}$ [26].
For parametric feedback cooling, the signal of each photodiode is sent to an FPGA board, which functions as a digital phase-locked loop (PLL), with the recovered signal filtered and phase-shifted to generate a feedback signal at twice the natural oscillation frequency $2\Omega_{q}$. These feedback signals are used to modulate the power of the trapping beam via an acousto-optic modulator (AOM).
In Fig. 2(a) we plot the power spectral densities (PSDs) of the signals measured at each output channel of the spatial mode sorter. We find that each output channel contains information on primarily a single translational DOF, equivalent to the nanoparticle’s motion along a particular spatial axis, encoded in $S_{qq}(\Omega)$ for $q=x,y$ and $z$. In particular, we observe good extinction ratios on the order of approximately $20$ dB, with the dominant spectral feature in each output channel being at least two orders of magnitude larger than all others. We plot these features, marking the shot noise floor for each output channel, in Fig. 2(b). These extinction ratios match the technical specification supplied for the spatial mode sorter for the LP11a and LP11b channels, suggesting that we may be limited by the performance capabilities of the device.
To quantify, and better understand the extent and limitations of this selectivity, we have numerically modelled the coupling efficiency of the scattered fields to each LP mode, defined as the spatial overlap integral
@@
where $\textbf{LP}_{\ell,m}(x,y)$ refers to the electric field of the LPℓ,m mode, and integration is performed over the end-face of the input optical fiber. This procedure, including the corrections made to account for the geometry of our setup, can be found in Appendix B (see Supplementary Material). Finally, note that we can write down analogous expressions for the coupling of the elastically scattered and LO fields also. The coefficients calculated for the particular experimental setup used here are reported in Fig. 3, with a detailed discussion on how we expect these quantities to relate to the measured PSDs found in Appendix D.
@@
Figure 3: The coupling efficiency of each electric field component to each LP mode supported by the spatial mode sorter, in the focal plane of the collections lens. We find that the inelastically scattered field components, $\bm{E}^{(q)}(x,y)$, selectively couple to specific LP modes. Note that the LP11a/b modes are effectively de-coupled from the elastically scattered and LO fields, $\bm{E}_{0}(x,y)$ and $\bm{E}_{\mathrm{ref}}(x,y)$, respectively.
The results shown in Fig. 3 qualitatively agree with the experimental results, with the LP11a (LP11b) mode coupling predominantly to the inelastically scattered field generated by displacements along $\bm{\hat{e}}_{x}$ ($\bm{\hat{e}}_{y}$), and the LP01 mode containing only weak contributions from both the transverse DOFs. Interestingly, the elastically scattered and LO fields are only weakly coupled to the LP11a/b modes, due to their spatial asymmetry. This would typically be disadvantageous in conventional detection systems, which rely on the LO to extract information from the phase of the scattered light. However, due to the coherent nature of the coupling described in Eq. (2), this is not the case in our approach. Indeed, any phase or amplitude modulation of the input field is naturally transformed to amplitude modulation of the LP modes upon coupling into the spatial mode sorter.
That is not to say that the decoupling of the LP11a/b modes from the LO field is inconsequential. Indeed, we find that in the absence of the LO field the measurement imprecision reduces to a much simpler form. This form is derived in Appendix C (see Supplementary Material) where we show that
$S^{\ell,m}_{\mathrm{imp},q}$ $=\frac{\hbar kc}{8\pi}\sqrt{\frac{\mu_{0}}{\varepsilon_{0}}}\frac{1}{\alpha_{q}^{2}}\frac{w_{0}^{2}}{(\eta^{\ell,m}_{q})^{2}}~{},$ (3)
where $\alpha_{0}$ and $\alpha_{q}$ are the real-valued amplitudes of the elastically and inelastically scattered electric field components, respectively. Counter-intuitively, this implies that the measurement imprecision is independent of the overlap of $\bm{E}^{(q)}(x,y)$ and the elastically scattered or LO fields, despite the interference of these fields being required to generate terms linearly proportional to the nanoparticle’s motion in the recovered power (see Appendix C in the Supplementary material). More importantly, it implies that, due to the large values of $\eta^{\ell,m}_{q}$ reported in Fig. 3, our technique should enable for higher measurement efficiencies than observed in conventional detection systems despite filtering out significant amounts of the incident light. This advantage is reflected in the signal-to-noise ratios observed in Fig. 2(b), which we note are unusually high for the transverse DOFs, with an optimal value of ${\mathrm{SNR}_{x}^{\mathrm{LP_{11a}}}\approx 110}$ dB above the shot noise floor.
@@
Finally, note that the LO field, as the dominant electric field component, usually also serves to offset the recovered signal, ensuring it sits above the system’s technical noise [27]. Since the LO does not couple into the LP11a/b modes to provide this offset, the power in these channels is much weaker than in the LP01 mode ($<10\%$). Consequently, if one were to measure every LP mode simultaneously with a single detector, $S_{zz}(\Omega)$ would dominate the measurement result, and the transverse DOFs would be difficult to discern. This is why a spatial mode sorter must be used, as opposed to a simple few-mode fiber configuration, and explains why the benefits of this approach have yet to be observed in previous experiments.
@@
As a proof-of-principle demonstration of our detection technique, we use parametric feedback cooling to control the COM motion of a levitated nanoparticle. The temperature $T_{q}$ achieved along the $\hat{\bm{e}}_{q}$ axis is determined by fitting the PSD measured at the appropriate output channel of the photonic chip (see Supplementary Material, Appendix H). To capture the actual dynamics of the system, the collected signals have been averaged over many periods of oscillation. Using this setup, we cool the motion of the nanoparticle from thermal equilibrium ($300$ K) to a minimum temperature of ${(T_{x},T_{y},T_{z})=(7,106,21.4)\pm(3.1,12,2.3)}$ mK, respectively. Interestingly, the minimum temperatures achieved here are an order of magnitude below what has been reported in the literature at pressures of approximately $5\times 10^{-5}$ mbar [21], the current technical limitation of our vacuum chamber. Indeed, we find that our observed values are approaching the theoretically predicted minimum values at this pressure (see Supplementary Material, Appendix I).
@@
We note that $T_{x}$ is the lowest observed temperature despite having the lowest reported measurement efficiency. We believe this is because $\Omega_{y}\approx 2\Omega_{z}$, such that the feedback signal generated for $z(t)$ inadvertently drives the nanoparticle’s motion along $\hat{\bm{e}}_{y}$. This can be possibly avoided by varying the ellipticity of the polarisation of the trapping beam, which tunes the transverse oscillation frequencies [28]. However, this also leads to modifications in the spatial profile of the scattered field, to the detriment of the demultiplexing protocol (see Supplementary Material, Appendix J).
@@
Insertion Losses $0.79$
Polarization-dependent Losses $0.99$
Coupling Efficiency $\begin{cases}0.71&,~{}x(t)\to\mathrm{LP}_{11\mathrm{a}}\\ 0.81&,~{}y(t)\to\mathrm{LP}_{11\mathrm{b}}\\ 0.92&,~{}z(t)\to\mathrm{LP}_{01}\\ \end{cases}$
@@
Measurement Efficiency $\left(\eta^{(q)}_{\mathrm{tot}}\right)$$\begin{cases}0.14&,~{}x(t)\to\mathrm{LP}_{11\mathrm{a}}\\ 0.16&,~{}y(t)\to\mathrm{LP}_{11\mathrm{b}}\\ 0.32&,~{}z(t)\to\mathrm{LP}_{01}\end{cases}$
@@
Table 1: Total efficiencies for the setup considered here. For a discussion on how each term has been determined, refer to the Supplementary Material, Appendix E.
In an effort to explain the improved performance of our parametric feedback cooling, we have opted to calculate the measurement efficiencies of this setup for each translational DOF. To understand why this quantity is of interest, note that optical measurements necessarily disturb the system’s natural evolution due to back-action. However, efficient measurements maximise the information gained per disturbance, and so can generate real-time feedback which not only controls the system’s mechanical motion but also cancels the effects of said back-action.
@@
We outline the information efficiencies of our system in Table 1, where we report record measurement efficiencies for the transverse DOFs. Importantly, this is not at the expense of the longitudinal DOF, for which we achieve a value comparable to previous experimental demonstrations of 1D ground state cooling [3, 2]. From these values, we can estimate the minimum achievable phonon occupancy for the nanoparticle’s COM motion along each spatial axis [29, 30, 3] in the limit of low pressures, where environmental information loss from random collisions with background gas molecules is negligible (${\approx 1\times 10^{-8}}$ mbar) [31]. Under these conditions, with an optimised feedback circuit, we expect to reach phonon occupancies of ${\bar{n}_{\mathrm{min}}^{q}=(1~{}/~{}\sqrt{\eta^{{}_{q}}_{{}^{\mathrm{tot}}}}-1)~{}/~{}2}$. For the setup reported here, this is equivalent to ${(\bar{n}_{\mathrm{min}}^{x},\bar{n}_{\mathrm{min}}^{y},\bar{n}_{\mathrm{min}}^{z})=(0.84,0.75,0.38)}$ phonons. Hence, this is the first demonstration of a detection system capable of realizing the 3D motional quantum ground state of a levitated nanoparticle using measurement-based control techniques.
@@
In summary, we have realized a novel approach to measuring the center-of-mass motion of a levitated dipolar scatterer, wherein the scattered field couples into an on-chip spatial mode sorter. By decomposing the inelastically scattered contributions to this field in the LP mode basis of an optical fiber, we can isolate contributions from each translational DOF in the amplitude of orthogonal optical modes. Thanks to this technique, we report record high, quantum-enabling measurement efficiencies for the translational degrees of freedom, which suggests that our technique can realize the 3D motional quantum ground state. Importantly, we believe that our technique can be extended to higher-order modes in the LP basis and generalized to other families such as Hermite-Gaussian (HG) modes, orbital angular momentum (OAM) modes, and Laguerre-Gaussian (LG) modes. This opens an exciting avenue for tracking the entire 6D (rotational and translational) motion of levitated objects with arbitrarily high precision. Alternatively, one could harness the toolbox of structured light to match the inelastically scattered field of the levitated object perfectly.
It is worth acknowledging that although the present demonstration focuses on the field of levitated optomechanics, the technique has the potential to be employed with other trapped microscopic objects such as trapped ions or neutral atoms where independently and directly accessing their different motional states, i.e. bosonic modes, could prove practical for hybrid qubit-oscillator quantum information processing [32, 33, 34].
Finally, the ability of the spatial mode sorter to isolate the scattered electric field associated with particular motional degrees of freedom is an essential step towards the passive, all-optical coherent feedback control of a levitated particle in free space. Close to the ground state, coherent feedback has the potential to improve cooling by avoiding measurement and feedback noise [19] and provides an avenue to generate mechanical squeezing [35].